OpenStudy (anonymous):

Solve the equation for x: 3x2 = 6x

OpenStudy (lgbasallote):

isolate x... 3x^2-6x = 0 you can use factoring here.. 3x(x-2) = 0 3x = 0 x - 2 =0 so x = 0 x=2 yes?

OpenStudy (anonymous):

I thought x is 1? 3x2 is 6. 6 /6 is 1.